# When Reasoning Meets Compression: Benchmarking Compressed Large Reasoning Models on Complex Reasoning Tasks

### v1

@@ 5 Results and Analysis
R1-Distill-Llama 70B Distillation 63.3 78.8 100.0 80.7 (13.0, 21.80)
R1-Distill-Llama 70B Distillation & 50% sparse 26.7 70.9 97.2 64.9 (6.0, 12.75)
R1-Distill-Qwen 32B Distillation 66.7 82.3 100.0 83.0 (1.0, 9.38)
R1-Distill-Qwen 32B Distillation & 50% sparse 30.0 75.4 96.0 67.1 (3.0, 9.29)
Table 2: Overall results of R1 and it compressed variants. The avg column represents the average accuracy scores of AIME 2024, FOLIO, and Temporal. EM and F1 of MuSiQue are shown in tuples. We segment this table based on model family (e.g., R1-Distill-Llama).
5.1 Overall Performance
The overall results of R1 and its compressed variants are presented in Table 2 and 3. We only show performance of pruned R1-Distill-Llama and pruned R1-Distill-Qwen under 50% sparsity in Table 2, as it is the default sparsity level of SparseGPT and many other works (Zhang et al., 2024a; Sun et al., 2023). Table 3 displays performance of our two distilled models under various sparsity levels.
Comparing Performance of Compression Methodologies In Table 2, 2.51-bit R1 achieves the highest average accuracy scores, surpassing even the original R1. We suspect that certain security mechanisms in DeepSeek may have lowered R1’s performance, as we often receive empty responses when using its API, requiring us to prompt R1 iteratively until valid responses are obtained. Both R1-Distill-Llama and R1-Distill-Qwen have competitive average accuracy scores. According to Table 3, we see that pruning also maintains good average accuracy scores when sparsity levels are low (e.g., 30% sparsity of R1-Distill-Llama). On MuSiQue dataset, the 2.51-bit model has the same EM as R1, with a slight decrease in F1. According to scores of the closed-book setting on GPT-4o (Zhang et al., 2025), all quantized models achieve relatively strong performance on MuSiQue, with higher EM and F1 than distilled and pruned models. Therefore, this 2.51-bit model has the best overall performance than other compression methods.
@@ 5.1 Overall Performance
Comparing Benchmark Difficulties We can make judgment about the difficulties of our accuracy-based benchmarks through Table 2. By comparing the scores of AIME 2024, FOLIO, and Temporal, we see all models struggle more on AIME 2024. This indicates that AIME 2024 is more difficult than the other two benchmarks. MuSiQue is also a difficult benchmark in terms of knowledge requirement, because its scores in Table 2 are much lower than RAG setup (Zhang et al., 2025; Jimenez Gutierrez et al., 2024). This suggests that existing LRMs lack sufficient knowledge for knowledge-intensive tasks, making RAG a more suitable approach. We demonstrate a diverse selection of reasoning benchmarks with varying difficulty levels, which is important for comprehensive evaluation.
Comparing Distilled Models based on LLaMA and Qwen When focusing on accuracy-based benchmarks, we see that R1-Distill-Qwen delivers an average 2.3% improvement over R1-Distill-Llama. This demonstrates that Qwen-32B has stronger reasoning capability than LLaMA-70B, which also aligns with the evaluation results of DeepSeek-R1 report (Guo et al., 2025). However, R1-Distill-Qwen scores significantly lower on MuSiQue, highlighting its inability to memorize detailed knowledge.
5.2 Collapse Point of Each Compression Method
We investigate whether LRMs degrade as they undergo increasing levels of compression. In Table 2, the performance of dynamically quantized models steadily declines as we move from 2.51-bit to 1.58-bit. However, we do not observe a clear collapse point, as performance declines only slightly when transitioning to a lower bit-size tier. As for distillation, comparing R1 with R1-Distill-Llama, we also do not observe a clear collapse point, although the performance drop on MuSiQue is significant. However, we see that R1-Distill-Qwen collapses on MuSiQue. With EM close to zero, it almost loses its ability on MuSiQue during distillation.
Comparing distilled models with their sparsified variants in Table 3, we find that pruned models collapse on all benchmarks at certain sparsity levels. Interestingly, their collapse point correlates to the difficulty of the benchmark. For example, on AIME 2024, R1-Distill-Llama collapses between 40% and 50% sparsity, since its performance drops by more than half. However, its collapse points on FOLIO and Temporal are roughly between 60% and 70% sparsity, which occur much later than AIME 2024. The correlation between collapse point and benchmark difficulty can also be seen on sparsified R1-Distill-Qwen. The early occurrence of collapse on AIME 2024 demonstrates that it is one of the most challenging datasets, as existing pruning works typically observe model collapse after 50% sparsity (Zhang et al., 2024a; Bai et al., 2024). Moreover, since R1-Distill-Qwen already collapses on MuSiQue, pruning it becomes meaningless for this dataset, as we observe very low scores for all sparsified R1-Distill-Qwen models on MuSiQue.
@@ 5.2 Collapse Point of Each Compression Method
Table 3: Performance of our two distilled models under various sparsity levels. The avg column represents the average accuracy scores of AIME 2024, FOLIO, and Temporal.

### v2

@@ 3.1 Overall Performance
R1-Distill-Llama 70B Distillation 65.6 79.8 99.9 81.8 (13.3, 21.57)
R1-Distill-Llama 70B Distillation & 50% sparse 23.3 71.6 97.6 64.2 (6.7, 13.49)
R1-Distill-Llama 70B Distillation & 4-bit AWQ 63.4 78.5 99.3 80.4 (10.7, 19.23)
R1-Distill-Llama 70B Distillation & 4-bit GPTQ 66.7 77.0 99.9 81.2 (10.3, 18.17)
R1-Distill-Llama 70B Distillation & 4-bit GPTAQ 64.4 78.8 99.6 80.9 (12.0, 21.57)
R1-Distill-Llama 70B Distillation & 3-bit GPTQ 46.7 71.8 99.3 72.6 (4.7, 11.92)
R1-Distill-Llama 70B Distillation & 3-bit GPTAQ 54.4 77.3 99.7 77.1 (5.7, 13.21)
R1-Distill-Qwen 32B Distillation 64.4 82.3 99.9 82.2 (2.7, 10.95)
R1-Distill-Qwen 32B Distillation & 50% sparse 25.6 75.1 97.9 66.2 (2.3, 9.01)
R1-Distill-Qwen 32B Distillation & 4-bit AWQ 67.8 82.3 99.1 83.1 (3.3, 10.28)
R1-Distill-Qwen 32B Distillation & 4-bit GPTQ 68.9 80.6 99.6 83.0 (4.0, 11.78)
R1-Distill-Qwen 32B Distillation & 4-bit GPTAQ 63.3 81.5 99.7 81.5 (2.7, 11.88)
R1-Distill-Qwen 32B Distillation & 4-bit ANY4 68.9 78.0 99.7 82.2 (5.7, 12.68)
R1-Distill-Qwen 32B Distillation & 3-bit GPTQ 44.4 74.2 98.9 72.5 (4.0, 11.55)
R1-Distill-Qwen 32B Distillation & 3-bit GPTAQ 45.6 77.5 99.5 74.2 (2.3, 9.18)
R1-Distill-Qwen 32B Distillation & 3-bit ANY3 53.3 82.6 99.9 78.6 (3.7, 10.27)
R1-Distill-Llama 8B Distillation 42.2 71.9 81.5 65.2 (0.0, 4.43)
R1-Distill-Llama 8B Distillation & 4-bit AWQ 47.8 68.0 84.0 66.6 (0.3, 5.05)
R1-Distill-Llama 8B Distillation & 4-bit GPTQ 42.2 66.2 65.9 58.1 (0.3, 4.68)
R1-Distill-Llama 8B Distillation & 4-bit GPTAQ 40.0 66.4 69.3 58.6 (0.0, 3.73)
R1-Distill-Llama 8B Distillation & 4-bit ANY4 41.1 68.5 88.7 66.1 (0.0, 3.54)
R1-Distill-Llama 8B Distillation & 3-bit GPTQ 11.1 65.0 67.3 47.8 (0.0, 2.89)
R1-Distill-Llama 8B Distillation & 3-bit GPTAQ 7.8 65.5 57.2 43.5 (0.0, 3.45)
R1-Distill-Llama 8B Distillation & 3-bit ANY3 3.3 50.1 34.9 29.4 (0.7, 2.35)
R1-Distill-Qwen 7B Distillation 46.7 78.0 75.6 66.8 (0.0, 3.57)
R1-Distill-Qwen 7B Distillation & 4-bit AWQ 46.6 75.5 74.9 65.7 (0.0, 3.14)
R1-Distill-Qwen 7B Distillation & 4-bit GPTQ 38.9 72.9 70.3 60.7 (1.0, 4.27)
R1-Distill-Qwen 7B Distillation & 4-bit GPTAQ 47.8 74.4 67.7 63.3 (0.0, 3.96)
R1-Distill-Qwen 7B Distillation & 4-bit ANY4 47.8 75.6 77.1 66.8 (0.0, 3.05)
R1-Distill-Qwen 7B Distillation & 3-bit GPTQ 17.8 65.7 31.7 38.4 (0.0, 3.12)
R1-Distill-Qwen 7B Distillation & 3-bit GPTAQ 24.4 64.5 48.7 45.9 (0.0, 3.06)
R1-Distill-Qwen 7B Distillation & 3-bit ANY3 32.2 69.3 30.1 43.9 (0.0, 3.89)
Comparing Compression Strategies. In Table 1, the 2.51-bit R1 achieves the highest average accuracy overall, since it has the smallest compression ratio. Both R1 distilled Llama-70B and Qwen-32B reach close-to-R1 accuracy scores. On MuSiQue, the 2.51-bit R1 also achieves performance close to original R1. Therefore, 2.51-bit R1 has the best overall performance than other compression strategies. Although R1 may be over-parameterized, a compression method with a smaller ratio can still offer advantages over methods with higher compression ratios. In contrast, pruning only 50% of the weights causes significant degradation, rendering the pruned LRMs unusable. Thus, we do not interpret the effect of pruning in the later sections. As for all distillation-only models, Qwen delivers stronger reasoning performance than Llama (Appendix E).
@@
Comparing Benchmark Difficulties. Comparing the scores using R1 distilled Llama-70B on AIME 2024, FOLIO, and Temporal, we see the largest score decrease on AIME 2024. This indicates that AIME 2024 is more challenging than the other two accuracy-based benchmarks. MuSiQue is also difficult in terms of knowledge requirement, because its scores in Table 1 are much lower than RAG (retrieval-augmented generation) setup (Zhang et al., 2025). This suggests that existing LRMs lack sufficient knowledge for knowledge-intensive tasks, making RAG a more suitable approach.
Takeaway 3.1 for Overall Performance Considering over-parameterization, methods with smaller compression ratios can still offer advantages over those with higher compression ratios. Regardless of whether compression is applied, LRMs lack sufficient knowledge for knowledge-intensive tasks.
3.2 Collapse Point
We investigate whether LRMs degrade as they undergo increasing levels of compression. In Table 1, the performance of dynamically quantized LRMs steadily declines as we move from 2.51 to 1.58-bit, but we do not observe a clear collapse point. All 4-bit AWQ, GPTQ, GPTAQ, and ANY4 reach performance similar to their unquantized counterparts, which shows the effectiveness of existing 4-bit quantization on LRMs. However, 3-bit GPTQ, GPTAQ, and ANY3 display signs of collapse, indicating bottlenecks of current 3-bit quantization. GPTAQ and ANY4 are newer than AWQ and GPTQ, and they achieve similar performance on 4-bit LRMs. Based on average accuracy, it is noteworthy that GPTAQ surpasses GPTQ on 3-bit LRMs for three out of four distilled models. Regarding distillation, R1 distilled Llama-8B and Qwen-7B achieve the lowest accuracy among all distillation-only models. Only R1 distilled Llama-70B yields decent MuSiQue scores.
Table 2: Performance of two distilled models under various sparsity levels. We report the one-pass scores for all models in this table.
@@ 3.2 Collapse Point
Table 2 displays performance of our two distilled models under various sparsity levels. Comparing distilled models with their sparsified variants, we find the precise collapse points of our pruned LRMs. Interestingly, their collapse points correlate to the benchmark difficulty. For example, on AIME 2024, R1-Distill-Llama collapses between 40% and 50% sparsity, since its performance drops by more than half. However, its collapse points on FOLIO and Temporal are roughly between 60% and 70% sparsity, which occur much later than AIME 2024. The correlation between collapse point and benchmark difficulty can also be seen on the sparsified Qwen.
Takeaway 3.2 for Collapse Point Collapse point correlates with benchmark difficulty. On hard benchmarks, 3-bit quantization and pruning with 50% sparsity or higher still have substantial room for improvement.
3.3 Compression Impact on Knowledge and Reasoning
In Table 2, although Qwen demonstrates stronger reasoning capabilities than Llama, it has significantly lower EM and F1 scores on MuSiQue. Because MuSiQue requires knowledge memorization under the closed-book setting, the smaller parameter count of Qwen puts itself at a disadvantaged position. In other words, models’ parameter count affects knowledge more than reasoning. In addition, we notice pruned R1-Distill-Llama-70B collapses between 30% and 40% sparsity on MuSiQue, which is even earlier than on AIME 2024. This shows pruning hurts LRMs’ knowledge memorization more than quantization. When a compression method aggressively removes the weights of an LRM, it is expected that the model’s knowledge will be more seriously affected. This phenomenon can also be seen on our dynamically quantized models in Table 1. Since quantization preserves parameter count and our analysis above shows that many quantized models still retain competitive reasoning capability, quantization is recommended on knowledge-intensive tasks.
Takeaway 3.3 for Compression Impact on Knowledge and Reasoning Pruning and distillation compress knowledge retention more than reasoning capabilities.
4 Distillation Effect on Weights
To study the effect of distillation on weights, we compute $\mathbf{I}_{m\ell}^{c}$ of two distilled R1 models and further measure the change of $\mathbf{RI}_{m\ell}^{c}$ as discussed in Section 2.3.
4.1 Locating Important Weights
The left part of Figure 2 presents the weight importance of R1 distilled Llama-8B in four heatmaps, each corresponding to a reasoning behavior. We observe that the final layer houses several most important linear modules across all four behaviors, with the highest value located at up_proj. Therefore, the up_proj in the final layer (32_up) stands out as the most important component.
Interestingly, this finding generalizes to R1 distilled Qwen-7B, as we also observe this up_proj outlier in the final layer of Qwen in Figure 4. Notably, our finding complements a recent analysis (Shao & Wu, 2025), which claims the most important module for reasoning is o_proj. Since identifying important weights is a core research problem of compression methodologies, our finding is valuable for future works.
Takeaway 4.1 for Locating Important Weights Distillation makes up_proj in the final layer as the most important module for reasoning behaviors, as observed in both R1 distilled Llama and Qwen models.
Figure 2: $\mathbf{I}_{m\ell}^{c}$ of DeepSeek-R1-Distill-Llama-8B (left) and change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to Llama-3.1-8B (right). Each heatmap displays scores of importance (or importance shift) of every module at each layer, providing a fine-grained analysis of weight contributions to the corresponding reasoning capability. On the right, increases in $\mathbf{RI}_{m\ell}^{c}$ are set to 0, as they only offset decreases elsewhere as discussed in 2.3. Every cluster of 4 side-by-side heatmaps (including those displayed below) follow the same scaling to show the precise magnitude of each weight module.
Figure 3: Change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to its 4-bit AWQ variant.
Table 3: Accuracy after selectively quantizing a single component of R1 distilled Llama-8B (e.g., 1_up means to only quantize the up_proj in the first layer) to 3-bit. Ranking of a component is based its $\sum_{c}\mathbf{I}_{m\ell}$, so “2nd col” refers to second place within its column across all four heatmaps (each column consists of 7 linear modules of a layer). “1st overall” means the global highest ranking.
Quantized Component Rank AIME 2024 FOLIO Temporal Avg
32_up 1st overall 20.0 63.1 63.6 48.9
32_gate 2nd col 33.3 62.1 67.2 54.2
32_v last col 43.3 68.0 79.6 63.6
31_up 2nd row 33.3 70.0 64.4 55.9
1_up last row 6.7 64.5 80.4 50.5
4.2 Validating Importance Scores
We validate Section 4.1 by applying 3-bit round-to-nearest quantization to either 32_up or a component sharing its column or row in the heatmaps, then measuring the resulting accuracy drop (Table 3). The more important a component is, the greater the accuracy drop when it is quantized. Specifically, we select four additional component candidates: the second- and last-ranked modules among the seven in the final layer, and the second- and last-ranked layers across all 32 layers of the up-projection. We see that 32_up yields the lowest average accuracy, which clearly demonstrates the validity of our findings in Section 4.1. It is quite salient that quantizing only this matrix (merely 0.7% of all weights) reduces the average accuracy by 16.3%. The component rank generally correlates with the accuracy drop, except for 1_up, which incurs the lowest accuracy on AIME 2024.
4.3 Importance Shift via Distillation
Since R1-Distill-Llama-8B is fine-tuned based on Llama-3.1-8B, we compute the change of $\mathbf{RI}_{m\ell}^{c}$ to visualize distillation effect in the right part of Figure 2 (for Llama) and 5 (for Qwen). Both parts of Figure 2 exhibit similar patterns (e.g., most outliers are in the final layer), indicating that the important weights of the distilled model are primarily the result of distillation with SFT, while the original Llama’s weight values play little role in shaping its reasoning capabilities. Thus, distillation effect is quite powerful in transforming a non-reasoning LLM into an LRM. For Qwen, Figures 4 and 5 also show similar patterns, so the utility of distillation generalizes to Qwen as well.
Takeaway 4.3 for Importance Shift via Distillation Important weights of the R1 distilled models are mainly the result of the distillation effect.
5 Quantization Effect on Weights
For quantization effect on weights, we analyze the decrease of importance shift during quantization.
5.1 Locating Quantization Effect
We show heatmaps to visualize the importance shift from R1 distilled Llama-8B to its 4-bit AWQ quantized variant in Figure 3. Across all four heatmaps, we observe a reduction in the significance of the gate projections in the middle layers (e.g., layer 9 to 23), suggesting that AWQ may overly compress these modules. Moreover, most linear modules in the final layer are compressed to the greatest extent, which shows the drawback of AWQ. Since 32_up is the most important module as discussed in Section 4.2 and its importance shift is little for uncertainty estimation and example testing capabilities, AWQ successfully preserves its significance on these two behaviors. However, for backtracking and adding knowledge capabilities, AWQ is not effective at maintaining its importance.
In Figure 6, we visualize the importance shift from R1 distilled Qwen-7B to its 4-bit AWQ quantized version. We also see a shift in the importance of the gate projections on Qwen, but this shift mainly occurs in the early layers (e.g., layer 1 to 10). On Qwen, AWQ does not preserve the importance of 32_up across all four reasoning capabilities, and it also overly compresses 32_k on two capabilities.
As another popular method, we interpret the effect of 4-bit GPTQ in Figure 7. On R1 distilled Llama-8B, we observe similar quantization effect as AWQ, since GPTQ also overly compresses final-layer modules and the gate projections in the middle layers. Their commonality demonstrates the generality of the bottlenecks we identified in existing quantization methods.
Takeaway 5 for Quantization Effect on Weights State-of-the-art quantization methods fail to preserve the importance of the MLP gate projections and the final layer, which is a key bottleneck of performance improvement.
5.2 Validating Quantization Effect
Table 4: Performance of 3-bit AWQ and selectively protecting the MLP modules in the final layer.
Model Compression Full-Precision Anywhere? AIME 2024 FOLIO Temporal Avg MuSiQue
R1-Distill-Llama-8B 3-bit AWQ - 10.0 59.6 68.4 46.0 (0.0, 3.50)
R1-Distill-Llama-8B 3-bit AWQ Final-layer MLP 16.7 67.0 74.0 52.57 (1.0, 3.62)
To validate our findings about the bottleneck of current quantization, we design a simple protection mechanism using a mixed precision fashion. We run two versions of 3-bit AWQ in Table 4. In the first version, we run AWQ with their default calibration data. Since we know AWQ overly compresses the MLP modules in the final layer, we then choose to protect them by changing their quantized weights to their original values in 16-bit. If they are truly important yet not well protected by AWQ, our protection mechanism should offer a significant improvement. Based on the discussion in Sections 3.1 and 3.2, we perform 3-bit quantization, since it will be the focus of future works.
We see our selective protection boosts 3-bit AWQ on all benchmarks, with an average accuracy improvement of 6.57%. This is particularly significant given that only about 2% of all weights remain in 16-bit. This mixed precision model outperforms all 3-bit quantization baselines in Table 1 by at least 4.77% in average accuracy, with gains of up to 23.17%. Therefore, our findings are demonstrated with an indication of substantial room for further improvement. Note that our protection provides relatively marginal increase on MuSiQue, as the weight count stays the same (Section 3.3).
6 Conclusion and Future Directions
We study the effects of LLMs compression on LRMs and present key findings for further imrpving LRMs compression. Future compression works are encouraged to consider the protection of MLP up projection in the final layer. The excessive compression of current quantization methods on MLP gate projections and final-layer modules highlights the need for preserving these weight modules.
Appendix A Use of LLMs
To improve the overall clarity of our writing, we used ChatGPT-4o and ChatGPT-5 via OpenAI’s web interface to polish a small fraction of sentences. LLMs were not used in any steps of the research ideation process. To ensure correctness and precision, we carefully reviewed and adapted all LLM-generated content before incorporating it into our writing.
Appendix B Related Work
Our literature review is conducted over existing compression methodologies (quantization, distillation, and pruning) and recent LRMs.
B.1 Quantization
PTQ is more popular in terms of the number of recent publications, because there is no retraining involved. For example, GPTQ (Frantar et al., 2022) and GPTAQ (Li et al., 2025) are one-shot weight quantization methods that use approximate second-order information, while AWQ (Lin et al., 2024) leverages activation distribution for finding the salient weight channels to skip. Other PTQ methods include weight-activation quantization (Shao et al., 2024; Yao et al., 2022; Liu et al., 2023) and KV cache quantization (Hooper et al., 2024; Liu et al., 2024b).
B.2 Distillation
Distillation involves two settings: black-box and white-box settings. For black-box setting, the teacher model is typically a closed-source LLM and only the outputs of teacher are available for the student model. For white-box setting, both weights and output distribution of the teacher model are available. Existing black-box distillation (Huang et al., 2024; Li et al., 2024b; Ho et al., 2023; Huang et al., 2022; Li et al., 2024a) prompts the teacher model to generate a training dataset for the student to learn. Specifically, researchers have started to distill OpenAI’s O1 model (Huang et al., 2024), which marks the beginning of LRMs compression. White-box distillation allows the student model to learn from teacher’s knowledge representation. Works has been done to align the output distribution (Agarwal et al., 2024; Gu et al., 2024) or the hidden representation (Liang et al., 2023) between teacher and student models.
B.3 Pruning
B.4 LRMs
Table 5: Dataset statistics of selected reasoning benchmarks.
Size Answer Type Metric Knowledge
AIME 2024 30 Integer Accuracy False
FOLIO 203 True/False/Uncertain Accuracy False
Temporal 250 (A)/(B)/(C)/(D) Accuracy False
MuSiQue 100 A few words (EM, F1) True
Appendix C Additional Details of Reasoning Benchmarks
Table 5 shows the statistics of our selected benchmarks. AIME 2024333https://huggingface.co/datasets/Maxwell-Jia/AIME_2024 (parts I and II) represents top match challenges, and its answers are integers. FOLIO [^4] https://huggingface.co/datasets/yale-nlp/FOLIO requires logical deductions to determine whether the provided conclusion is true, false, or uncertain based on premise. In Temporal Sequences [^5] https://github.com/suzgunmirac/BIG-Bench-Hard/blob/main/bbh/temporal_sequences.json, models are asked to use a provided timeline to determine what time a person might be free to perform another activity. Since each of its questions comes with four options, we expect our models to output the index (the letter) of the selected option. Since MuSiQue involves question answering and its answers are in a few words, we adopt exact match (EM) and F1. We randomly sample 100 questions out of 1000 from MuSiQue for our benchmarking analysis.
Appendix D Implementation Details
We run the dynamically quantized models on llama.cpp [^6] https://github.com/ggml-org/llama.cpp based on their requirement. We run all other distilled, pruned, and quantized models on vLLM (Kwon et al., 2023) for its fast inference. In order to comprehensively analyze performance change after compression, we also evaluate R1 on our reasoning benchmarks by using DeepSeek API. Aligning with DeepSeek-R1 report (Guo et al., 2025), we keep the same parameters for all models during inference: maximum generation length is set to $32768$, temperature is set to $0.6$, and top-p value is set to $0.95$.
We use AutoAWQ [^7] https://github.com/casper-hansen/AutoAWQ as the AWQ implementation for inference due to its speed advantage (vLLM support), while the original AWQ code [^8] https://github.com/mit-han-lab/llm-awq is used to generate the pseudo-quantized R1 distilled Llama-8B for our analysis in Section 5.
We focus on analyzing the effect of compression methods on performance and thus do not consider inference speedup. The reason is that these methods run on different inference platforms, so it is hard to control the consistency of inference optimization across various platforms.
Appendix E Comparing Distilled Models
On accuracy-based benchmarks of Table 1, we see that R1 distilled Qwen-32B delivers an average 0.4% improvement over Llama-70 and R1 distilled Qwen-7B delivers an average 1.6% improvement over Llama-8B. Although these two Qwen models have less weights, Qwen delivers stronger reasoning performance than Llama. This phenomenon aligns with DeepSeek report (Guo et al., 2025). However, R1-Distill-Qwen-32B scores significantly lower than R1-Distill-Llama-70B on MuSiQue, highlighting its worse ability of memorization.
Appendix F Additional Visualization of Weight Importance and Importance Shift
All additional figures in Appendix are thoroughly discussed in the main content. Figure 4 shows the weight importance of DeepSeek-R1-Distill-Qwen-7B across four heatmaps, each corresponding to a specific target reasoning behavior. Figure 5 displays the change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to Qwen2.5-Math-7B. To decode the quantization effect on Qwen, Figure 6 shows the change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Qwen-7B to its 4-bit AWQ variant. Similarly, Figure 7 shows the change of $\mathbf{RI}_{m\ell}^{c}$ from R1 distilled Llama-8B to its 4-bit GPTQ quantized variant.
Figure 4: $\mathbf{I}_{m\ell}^{c}$ of DeepSeek-R1-Distill-Qwen-7B.
Figure 5: Change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to Qwen2.5-Math-7B (the backbone model).
Figure 6: Change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Qwen-7B to its 4-bit AWQ variant.
Figure 7: Change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to its 4-bit GPTQ quantized variant.
Appendix G Test-time Compute
We study the behavior of R1 and some of its compressed variants by measuring their test-time compute. Table 6 shows the analysis of test-time compute when we select the shortest and longest 30% of responses output by each model on each benchmark. We observe that shorter model outputs consistently yield better performance across three reasoning benchmarks, with only an exception on MuSiQue. Regardless of whether an LRM is compressed, if it generates significantly more tokens for a question than other questions in the same dataset, the answer is likely to be incorrect. We exclude Temporal here, because many of the compressed R1 models achieve close to 100% accuracy. The length ratios between the longest and the shortest 30% are typically greater than 4, which indicates nontrivial length differences among model outputs.
Similarly, Table 7 shows the selection of the shortest and the longest 20% of model responses. Compared to Table 6, as we move toward the extreme, the performance gap between the shortest and longest responses becomes even larger. Both R1 and its compressed variants achieve higher scores when they spend less compute during test time. After manually checking long responses, we notice that longer outputs tend to be more verbose and involve more backtracking in reasoning. This finding demonstrates the need to reduce verbosity to improve reasoning performance, which aligns with recent research (Zhang et al., 2024b).